# VProChart: Answering Chart Question through Visual Perception Alignment Agent and Programmatic Solution Reasoning

### v1

@@ Abstract
Charts are widely used for data visualization across various fields, including education, research, and business. Chart Question Answering (CQA) is an emerging task focused on the automatic interpretation and reasoning of data presented in charts. However, chart images are inherently difficult to interpret, and chart-related questions often involve complex logical and numerical reasoning, which hinders the performance of existing models. This paper introduces VProChart, a novel framework designed to address these challenges in CQA by integrating a lightweight Visual Perception Alignment Agent (VPAgent) and a Programmatic Solution Reasoning approach. VPAgent aligns and models chart elements based on principles of human visual perception, enhancing the understanding of chart context. The Programmatic Solution Reasoning approach leverages large language models (LLMs) to transform natural language reasoning questions into structured solution programs, facilitating precise numerical and logical reasoning. Extensive experiments on benchmark datasets such as ChartQA and PlotQA demonstrate that VProChart significantly outperforms existing methods, highlighting its capability in understanding and reasoning with charts.
@@ 1 Introduction
Chart is a widely used format for data visualization, prevalent in scientific papers and business reports. Chart Question Answering (CQA) aims to answer questions based on chart contexts, achieving the automatic analysis of data trends and the automatic generation of data reports. In recent years, the CQA task has garnered increasing attention in light of advancements in multi-modal understanding and reasoning techniques. [9, 25]
@@
The CQA task naturally draws comparisons with Visual Question Answering (VQA), which involves answering questions based on images. Despite significant advances in VQA [15, 27], CQA remains challenging due to the complex understanding of diagram contexts and complex reasoning requirements. Firstly, chart images are inherently difficult to interpret. For example, in the left chart of Figure 1, the line chart does not have a strict legend, it only provides two labels corresponding to colors, which might potentially cause confusion. The elements and organizational methods in charts are not derived from nature but depend on the author’s intentions. Different authors may use various elements or organizational methods for clarity or aesthetic appeal, complicating the understanding of chart images through common presentation rules. Secondly, questions in CQA often require complex numerical or logical reasoning. For example, in the left chart of Figure 1, “Is the sum of the heights of the two highest red line greater than the value of the two highest blue line?” necessitates numerical calculations and logical comparisons. These types of reasoning are particularly challenging for current research methods. However, existing approaches lack significant improvement in both of these aforementioned aspects.
On the one hand, the currently dominant strategy includes models like Pix2Struct [13] and UniChart [20], which are end-to-end visual models, using extensive synthetic chart pretraining to enhance their chart understanding capabilities. This method has made progress in understanding synthetic charts but is still limited by the performance of external tools and synthetic data. On the other hand, MatCha [16] and ChartInstruct [22] have successfully enhanced the reasoning capabilities of models on CQA tasks through logical reasoning computation datasets or instruction fine-tuning. However, these reasoning approaches are still based on neural networks, making it difficult to achieve further progress in numerical reasoning. Overall, existing research lacks improvements for the two aforementioned challenges: in chart understanding, current methods rely on pretraining with large-scale chart-question pairs, while ignoring explicit modeling of relationships among chart elements; in numerical and logical reasoning, they rely on pretrained neural networks, which are not adept at handling complex logical and numerical reasoning.
@@
  • We conduct extensive experiments on different datasets such as ChartQA and PlotQA to validate the superiority of VProChart. The results show significant improvements compared to various popular competitors across different scales. We have made VProChart publicly available at https://github.com/MuyeHuang/VProChart
@@ 2 Related Work
The chart question answering task answers textual questions by understanding visual context. Kahou et al. [11] pioneered a synthetic Chart dataset FigureQA, which includes automatically generated chart images and a small number of template-generated true/false judgment questions. Kafle et al. [9] proposed another more comprehensive synthetic dataset, DVQA. They also introduced two models named MOM and SANDY to better handle CQA problems, and they subsequently introduced the PReFIL [10] model, achieving state-of-the-art results on FigureQA and DVQA.
However, the aforementioned datasets lack the complex reasoning problems present in real-world CQA. Therefore, PlotQA [25] introduced a large-scale dataset, containing complex reasoning and numerical computation tasks. It comprises real-world data along with diverse synthetic charts. Furthermore, Masry et al. [21] proposed the novel CQA dataset ChartQA, where the human subset consists of real-world images and manually curated questions, bringing the CQA task closer to real-world scenarios. CRCT [14] achieved satisfactory results on the PlotQA dataset by utilising a unified regression-classification head. Lee et al. [13] introduced a pure visual model Pix2Struct based on Vision Transformer [7], achieving state-of-the-art performance on various CQA datasets through large-scale pre-training on screenshots without the need for external OCR. Building upon this, MatCha [16] using mathematical reasoning tasks during pre-training, resulting in significant improvements on ChartQA and PlotQA. Masry et al. [20] proposed UniChart, which achieved state-of-the-art performance on the ChartQA task through multi-task multi-stage pre-training. In light of the rapid advancement of LLMs, ChartLlama [8] utilised Llama with an added visual branch, achieving good performance through large-scale pre-training and offering new directions for researchers. Moreover, ChartInstruct [22], ChartAst [24] and TinyChart [30], an instructive approach to the CQA domain, which achieved new advancements in CQA tasks through the utilisation of LLMs.
@@ 3.1 Visual Perception Alignment Agent
The encoding of chart images plays a critical role in the CQA process, as it requires the extraction of both visual and textual elements from the charts. To address the need for a lightweight and structurally complex encoder, we employ Swin-Encoder [19] as the chart encoder. Swin-Encoder is a lightweight encoder capable of achieving high-resolution encoding at a relatively low computational cost, enabling the simultaneous extraction of visual and textual information from images without the need for external OCR.
@@
The Question Encoder Module encodes questions into sequences and extracts a global representation of each question. We employ RoBERTa [18], renowned for its state-of-the-art performance in natural language processing tasks, as the encoder. Given a question $\mathbf{Q}$, this procedure represents the question as a sequence:
@@
For a given question $\mathbf{Q}$, the answer can be either a floating-point number or textual content. Previous works, such as T5 [26], have demonstrated that generative decoders can effectively handle both types of responses simultaneously. Consequently, we employ the mBART [17] decoder as the text decoder. The text prompt is provided as input to the decoder, which generates the answer based on the prompt context. The loss function for this process is defined as follows:
@@ 3.2 Programmatic Solution Reasoning
1) Prompting LLMs.
Prompting LLMs leverage the complex question understanding and code generation capabilities of LLMs to transform intricate textual questions into solution program [28]. LLMs do not need to compute or infer answers, instead, they express the process of solving the problem in the form of solution program [4]. In particular, we prompt the LLMs to act as programming experts, solving problems programmatically by invoking two kinds of operators: 1) $ASK(\cdot)$. This operator takes the complete question as input and returns the answer directly, used for questions that do not require reasoning or computation. For instance, the question “Which color indicates 65+ years?” requires no reasoning. 2) $SUBSTEP(\cdot)$. This operator takes in sub-questions and returns answers to these sub-questions, used for complex questions requiring decomposition.
2) Python-style Reasoning Engine.
4 Experiments
@@ 4.1 Experimental Setting
Datasets. We conducted extensive experiments on three prominent datasets: PlotQA [25], DVQA [9], and ChartQA [21]. DVQA is a fully synthetic dataset focused on chart-specific questions. PlotQA, a large synthetic dataset, comprises two subsets, D1 and D2, incorporating both real-world data and automatically generated charts, with numerous numerical and logical reasoning questions. ChartQA includes two benchmarks for challenging CQA tasks: ChartQA-H and ChartQA-M. ChartQA-H features manually curated chart images and question-answer pairs, while ChartQA-M consists of machine-generated images and questions. As illustrated in Figure 1, ChartQA-H poses a more formidable challenge due to its extensive collection of manually curated charts and complex multi-step reasoning questions.
Competitors. We evaluate our model against both fully-supervised models and large-language-model-based baselines. Fully-supervised models include T5 [26], PReFIL [10], VisionTapas [21], Pix2Struct [13], ChartReader [6], CRCT [14], MatCha [16], and UniChart [20], all of which are trained on CQA benchmarks. Large-language-model-based models such as PaLI-17B (res.588) [5], Qwen-VL [2], ChartLLama [8], OneChart [3], ChartGemma [23] and ChartInstruct [22] utilize end-to-end LLMs integrated with a visual component.
Settings. In our proposed VPAgent, we employ the weights of the Donut [12] for initialization. The training process consists of two stages: firstly, pre-training 200k steps on the UniChart pretraining corpus [20], with an initial learning rate of 1e-4; secondly, fine-tuning 5 epoch on the downstream dataset, with an initial learning rate of 5e-5. In our programmatic solution reasoning framework, we employ Qwen2-7B-Instruct [29] as the base LLM. All experiments were conducted using two Tesla A100 GPUs.
@@ 4.2 Performance Comparison
4.3 Analysis on VPAgent
VPAgent Ablation. To validate the contribution of the VP alignment module and the Q-driven reasoning module within VPAgent, we designed detailed ablation experiments for VPAgent, as shown in Table 3. We use “w/o” to indicate the absence of a particular module. It is noteworthy that since VP alignment is a prerequisite module for Q-driven chart reasoning, in the “w/o VP alignment” experiment, we substitute the alignment matrix $\mathbf{V}$ with Gaussian noise. We can make the following three observations:
@@ 4.3 Analysis on VPAgent
Method ChartQA
Human Augment Avg
Max Cluster Numbers
3-Clusters 44.2 87.1 65.7
4-Clusters 46.9 87.3 67.1
5-Clusters 48.1 87.4 67.8
Intra-Cluster Layers
3-Layers 44.7 88.6 66.7
2-Layers 46.6 88.0 67.3
1-Laters 48.1 87.4 67.8
Table 4: Architecture Study on VPAgent.
VPAgent Architecture Study. As shown in Table 4, we conducted ablation experiments on the number of clusters and the Intra-Cluster Reasoning layer within VPAgent. The settings employed by VPAgent are highlighted in gray in the table. In the max cluster numbers experiment, it can be observed that the performance of 3 and 4 clusters is significantly lower than that of 5 clusters. This is reasonable, as a more fine-grained clustering can enhance the applicability of the VP criterion. On the other hand, in the Intra-Cluster Layers experiment, 3 Layers outperform 1 Layer on the synthetic “Augmented” dataset, but underperform 1 Layer on the manually annotated “Human” dataset. This is an intriguing finding, demonstrating that manually annotated real-world scenarios do not necessitate deeper Intra-Cluster Reasoning Layers.
Agent Method RealCQA
Math Extract
Unichart 13.0 33.0
ChartLlama 10.0 13.0
MatCha 16.0 27.5
ChartAst 15.0 36.0
VPAgent 24.5 36.8
Table 5: Zero shot study on VPAgent
VPAgent Zero-Shot Study. To validate VPAgent performance on data unseen during training, we conducted a zero-shot experiment on the RealCQA [1] dataset. For evaluation purposes, we utilize accuracy with a 5% tolerance. Following the configuration of ChartAst [24], we treated the ‘4b’ category in RealCQA as the Math type and the ‘3a’ and ‘3b’ categories as Extract questions. As illustrated in Table 5, VPAgent demonstrates superior performance in zero-shot mathematical extraction (Math) and textual extraction (Extract) compared to models like Unichart and MatCha.
4.4 Analysis on Programmatic Solution
@@ 5 Conclusion
In this paper, we propose the VProChart, a framework comprising a VPAgent guided by human visual principles and Programmatic Solution Reasoning for precise inference, which presents a novel approach to the ChartQA task. We conducted comprehensive experiments on ChartQA, PlotQA, and DVQA, demonstrating the effectiveness of VProChart. In the future, we will continue to explore the potential of integrating human-like thinking with programmatic reasoning.
6 Appendix
6.1 Detialed Case Study
More case studies of VProChart are illustrated in Figures 5 and 6.
Figure 5: A Detailed Runing Example 1 of VProChart.
Figure 6: A Detailed Runing Example 2 of VProChart.

### v2

@@ Abstract
Charts are widely used for data visualization across various fields, including education, research, and business. Chart Question Answering (CQA) is an emerging task focused on the automatic interpretation and reasoning of data presented in charts. However, chart images are inherently difficult to interpret, and chart-related questions often involve complex logical and numerical reasoning, which hinders the performance of existing models. This paper introduces VProChart, a novel framework designed to address these challenges in CQA by integrating a lightweight Visual Perception Alignment Agent (VPAgent) and a Programmatic Solution Reasoning approach. VPAgent aligns and models chart elements based on principles of human visual perception, enhancing the understanding of chart context. The Programmatic Solution Reasoning approach leverages large language models (LLMs) to transform natural language reasoning questions into structured solution programs, facilitating precise numerical and logical reasoning. Extensive experiments on benchmark datasets such as ChartQA and PlotQA demonstrate that VProChart significantly outperforms existing methods, highlighting its ability to understand and reason with charts.
Homepagehttps://github.com/MuyeHuang/VproChart
@@ 1 Introduction
Chart is a widely used format for data visualization, prevalent in scientific papers and business reports. Chart Question Answering (CQA) aims to answer questions based on chart contexts, achieving the automatic analysis of data trends and the automatic generation of data reports. In recent years, the CQA task has garnered increasing attention in light of advancements in multi-modal understanding and reasoning techniques. (Kafle et al. 2018; Methani et al. 2020)
The CQA task naturally draws comparisons with Visual Question Answering (VQA), which involves answering questions based on images. Despite significant advances in VQA (Li et al. 2023; Wang et al. 2023), CQA remains challenging due to the complex understanding of diagram contexts and complex reasoning requirements. Firstly, chart images are inherently difficult to interpret. For example, in the left chart of Figure 1, the line chart does not have a strict legend, it only provides two labels corresponding to colors, which might potentially cause confusion. The elements and organizational methods in charts are not derived from nature but depend on the author’s intentions. Different authors may use various elements or organizational methods for clarity or aesthetic appeal, complicating the understanding of chart images through common presentation rules. Secondly, questions in CQA often require complex numerical or logical reasoning. For example, in the left chart of Figure 1, “Is the sum of the heights of the two highest red line greater than the value of the two highest blue line?” necessitates numerical calculations and logical comparisons. These types of reasoning are particularly challenging for current research methods. However, existing approaches lack significant improvement in both of these aforementioned aspects.
@@
On the one hand, the currently dominant strategy includes models like Pix2Struct (Lee et al. 2023) and UniChart (Masry et al. 2023), which are end-to-end visual models, using extensive synthetic chart pretraining to enhance their chart understanding capabilities. This method has made progress in understanding synthetic charts but is still limited by the performance of external tools and synthetic data. On the other hand, MatCha (Liu et al. 2023) and ChartInstruct (Masry et al. 2024a) have successfully enhanced the reasoning capabilities of models on CQA tasks through logical reasoning computation datasets or instruction fine-tuning. However, these reasoning approaches are still based on neural networks, making it difficult to achieve further progress in numerical reasoning. Overall, existing research lacks improvements for the two aforementioned challenges: in chart understanding, current methods rely on pretraining with large-scale chart-question pairs, while ignoring explicit modeling of relationships among chart elements; in numerical and logical reasoning, they rely on pretrained neural networks, which are not adept at handling complex logical and numerical reasoning.
@@
  • We conduct extensive experiments on different datasets such as ChartQA and PlotQA to validate the superiority of VProChart. The results show significant improvements compared to various popular competitors across different scales.
@@ 2 Related Work
The chart question answering task answers textual questions by understanding visual context. Kahou et al. (Kahou et al. 2018) pioneered a synthetic Chart dataset FigureQA, which includes automatically generated chart images and a small number of template-generated true/false judgment questions. Kafle et al. (Kafle et al. 2018) proposed another more comprehensive synthetic dataset, DVQA. They also introduced two models named MOM and SANDY to better handle CQA problems, and they subsequently introduced the PReFIL (Kafle et al. 2020) model, achieving state-of-the-art results on FigureQA and DVQA.
However, the aforementioned datasets lack the complex reasoning problems present in real-world CQA. Therefore, PlotQA (Methani et al. 2020) introduced a large-scale dataset, containing complex reasoning and numerical computation tasks. It comprises real-world data along with diverse synthetic charts. Furthermore, Masry et al. (Masry et al. 2022) proposed the novel CQA dataset ChartQA, where the human subset consists of real-world images and manually curated questions, bringing the CQA task closer to real-world scenarios. CRCT (Levy, Ben-Ari, and Lischinski 2022) achieved satisfactory results on the PlotQA dataset by utilising a unified regression-classification head. Lee et al. (Lee et al. 2023) introduced a pure visual model Pix2Struct based on Vision Transformer (Dosovitskiy et al. 2021), achieving state-of-the-art performance on various CQA datasets through large-scale pre-training on screenshots without the need for external OCR. Building upon this, MatCha (Liu et al. 2023) using mathematical reasoning tasks during pre-training, resulting in significant improvements on ChartQA and PlotQA. Masry et al. (Masry et al. 2023) proposed UniChart, which achieved state-of-the-art performance on the ChartQA task through multi-task multi-stage pre-training. In light of the rapid advancement of LLMs, ChartLlama (Han et al. 2023) utilised Llama with an added visual branch, achieving good performance through large-scale pre-training and offering new directions for researchers. Moreover, ChartInstruct (Masry et al. 2024a), ChartAst (Meng et al. 2024) and TinyChart (Zhang et al. 2024), an instructive approach to the CQA domain, which achieved new advancements in CQA tasks through the utilisation of LLMs.
@@ 3.1 Visual Perception Alignment Agent
The encoding of chart images plays a critical role in the CQA process, as it requires the extraction of both visual and textual elements from the charts. To address the need for a lightweight and structurally complex encoder, we employ Swin-Encoder (Liu et al. 2021) as the chart encoder. Swin-Encoder is a lightweight encoder capable of achieving high-resolution encoding at a relatively low computational cost, enabling the simultaneous extraction of visual and textual information from images without the need for external OCR.
@@
The Question Encoder Module encodes questions into sequences and extracts a global representation of each question. We employ RoBERTa (Liu et al. 2019), renowned for its state-of-the-art performance in natural language processing tasks, as the encoder. Given a question $\mathbf{Q}$, this procedure represents the question as a sequence:
@@
For a given question $\mathbf{Q}$, the answer can be either a floating-point number or textual content. Previous works, such as T5 (Raffel et al. 2020), have demonstrated that generative decoders can effectively handle both types of responses simultaneously. Consequently, we employ the mBART (Liu et al. 2020) decoder as the text decoder. The text prompt is provided as input to the decoder, which generates the answer based on the prompt context. The loss function for this process is defined as follows:
@@ 3.2 Programmatic Solution Reasoning
1) Prompting LLMs.
Prompting LLMs leverage the complex question understanding and code generation capabilities of LLMs to transform intricate textual questions into solution program (Xu et al. 2023). LLMs do not need to compute or infer answers, instead, they express the process of solving the problem in the form of solution program (Chen et al. 2023a). In particular, we prompt the LLMs to act as programming experts, solving problems programmatically by invoking two kinds of operators: 1) $ASK(\cdot)$. This operator takes the complete question as input and returns the answer directly, used for questions that do not require reasoning or computation. For instance, the question “Which color indicates 65+ years?” requires no reasoning. 2) $SUBSTEP(\cdot)$. This operator takes in sub-questions and returns answers to these sub-questions, used for complex questions requiring decomposition.
2) Python-style Reasoning Engine.
4 Experiments
@@ 4.1 Experimental Setting
Datasets. We conducted extensive experiments on three prominent datasets: PlotQA (Methani et al. 2020), DVQA (Kafle et al. 2018), and ChartQA (Masry et al. 2022). DVQA is a fully synthetic dataset focused on chart-specific questions. PlotQA, a large synthetic dataset, comprises two subsets, D1 and D2, incorporating both real-world data and automatically generated charts, with numerous numerical and logical reasoning questions. ChartQA includes two benchmarks for challenging CQA tasks: ChartQA-H and ChartQA-M. ChartQA-H features manually curated chart images and question-answer pairs, while ChartQA-M consists of machine-generated images and questions. As illustrated in Figure 1, ChartQA-H poses a more formidable challenge due to its extensive collection of manually curated charts and complex multi-step reasoning questions.
Competitors. We evaluate our model against both fully-supervised models and large-language-model-based baselines. Fully-supervised models include T5 (Raffel et al. 2020), PReFIL (Kafle et al. 2020), VisionTapas (Masry et al. 2022), Pix2Struct (Lee et al. 2023), ChartReader (Cheng, Dai, and Hauptmann 2023), CRCT (Levy, Ben-Ari, and Lischinski 2022), MatCha (Liu et al. 2023), and UniChart (Masry et al. 2023), all of which are trained on CQA benchmarks. Large-language-model-based models such as PaLI-17B (res.588) (Chen et al. 2023b), Qwen-VL (Bai et al. 2023), ChartLLama (Han et al. 2023), OneChart (Chen et al. 2024), ChartGemma (Masry et al. 2024b) and ChartInstruct (Masry et al. 2024a) utilize end-to-end LLMs integrated with a visual component.
Settings. In our proposed VPAgent, we use the Donut weights (Kim et al. 2022) for initialization. The training process consists of two stages: firstly, pre-training 200k steps on the UniChart pretraining corpus (Masry et al. 2023), with an initial learning rate of 1e-4; secondly, fine-tuning 5 epoch on the downstream dataset, with an initial learning rate of 5e-5. Notably, due to the substantial computational cost, we trained the model only on PlotQA-D1 and tested it on PlotQA-D2. In our programmatic solution reasoning framework, we employ Qwen2-7B-Instruct (Yang et al. 2024) as the base LLM. All experiments were conducted using two Tesla A100 GPUs.
@@ 4.2 Performance Comparison
4.3 Analysis on VPAgent
VPAgent Ablation. To validate the contribution of the VP alignment module and the Q-driven reasoning module within VPAgent, we designed detailed ablation experiments for VPAgent, as shown in Table 3. Further experiments, including the zeroshot study and architecture study, can be found in the appendix. We use “w/o” to indicate the absence of a particular module. It is noteworthy that since VP alignment is a prerequisite module for Q-driven chart reasoning, in the “w/o VP alignment” experiment, we substitute the alignment matrix $\mathbf{V}$ with Gaussian noise. We can make the following three observations:
@@ 5 Conclusion
In this paper, we propose the VProChart, a framework comprising a VPAgent guided by human visual principles and Programmatic Solution Reasoning for precise inference, which presents a novel approach to the ChartQA task. We conducted comprehensive experiments on ChartQA, PlotQA, and DVQA, demonstrating the effectiveness of VProChart, and showing that human visual perception can significantly enhance the model’s capability in understanding schematic diagrams. In the future, we will continue to explore the potential of integrating human-like thinking with programmatic reasoning.
Acknowledgments
This work was supported by National Key Research and Development Program of China (2022YFC3303600), the Key Research and Development Project in Shaanxi Province No. 2022GXLH-01-03, National Natural Science Foundation of China (62137002, 62192781, 62293553, 61937001, 62250066, 62306229, and 62106190), “LENOVO-XJTU” Intelligent Industry Joint Laboratory Project, Foundation of Key National Defense Science and Technology Laboratory (6142101210201), Natural Science Basic Research Program of Shaanxi (2023-JC-YB-593), the Youth Innovation Team of Shaanxi Universities, the Shaanxi Provincial Social Science Foundation Project (2024P041).